Dick and Barbara moved to Ipswich and purchased a home. They paid $ 1856 per year to insure their two cars. Dick said it cost $ 385 more per year to insure his car than it did to insure Barbaras car. How much did it cost to insure Dicks car and how much to insure Barbaras car?
Can you tell me if I set this up correctly? x + 385 = y

Answer by 303795(602)

x is the cost to insure Barabara's car and y is the amount to insure Dicks car.
You have set up the problem correctly.
Now the total to insure both cars is 1856 so x + y = 1856 and y = x + 385
So x + x + 385 = 1856
2x + 385 = 1856
That should be enough for you to finish the problem off.